OpenStudy (anonymous):

simplify: (sin^2x)/(sec^2x-1)

OpenStudy (anonymous):

help :(

OpenStudy (anonymous):

sin^2 + cos^2 = 1 dividing through by cos^2 tan^2 + 1 = sec^2 therefore sec^2 -1 = tan^2

OpenStudy (anonymous):

so the question becomes\[\frac{\sin^2(x)}{\tan^2(x)}\]

OpenStudy (anonymous):

\[= \sin^2(x) \times \frac{\cos^2(x)}{\sin^2(x)}\]

OpenStudy (anonymous):

\[= \cos^2(x)\]

OpenStudy (anonymous):

remember that \[\tan(x) = \frac{\sin(x)}{\cos(x)}\] then when divide by something it is the same as flipping that"something", and multiplying